The Company’s beta is 1.25 and its dividend growth rate is 14.75%, just yesterday, it paid a dividend of $1.75.  Today’s share price is $53.00.  Furthermore, you believe that the share price moves in accordance with the dividend constant growth model.  The economy wide risk free interest rate is 4.5% and the expected risk premium for the market portfolio is 9.5%.  You believe that the stock represents a good investment if the expected total return implied by the dividend constant growth model exceeds the required rate of return implied by the Capital Asset Pricing Model.  What is the required rate of return and expected rate of return for the stock?  Should you buy it; why or why not?
Answer #1

Hi,

Here beta = 1.25

rf = 4.5%

market risk premium = 9.5%

As per CAPM

required rate of return r= rf + beta*market risk premium

   = 4.5 + 1.25*9.5

   = 16.375%

Current Dividend D0 = $1.75

growth rate g = 14.75%

As per dividend discount model

Current Share price = D0*(1+g)/(r-g)

53 = 1.75*(1+14.75%)/(r-14.75%)

53*(r-14.75%) = 2.008125

53r - 7.8175 = 2.008125

r = 18.54%

Hence expected return from DDM is greater than required return from CAPM so you should buy the Share.

Thanks